A car traveling due east with an initial velocity of 10 m/s accelerates for 6 seconds at a constant rate of 4 m/s2. i.What is its velocity at the end of this time? ii.How far does it travel during this time?


1
Expert's answer
2021-05-30T13:28:24-0400

From the kinematic equaiton of the motion with constant acceleration, the velocity after time "t = 6s" is:


"v = v_0 + at"

where "v_0 = 10m\/s" is the initial velocity, "a = 4m\/s^2" is the acceleration. Thus, obtain:


"v = 10m/s + 4m/s^2\cdot 6s = 34m/s"

From another kinematic equation, the distance travelled in the same time is:


"d = v_0t + \dfrac{at^2}{2}\\
d = 10m/s\cdot 6s + \dfrac{4m/s^2\cdot (6s)^2}{2} =132m"

Answer. 34 m/s and 132 m.